OpenStudy (anonymous):

1. Renting a bicycle at a local Bike shop costs $30 for the first day and $10 for each additional day. then (a) Create a chart of values and write an equation to find the total cost, C in terms of the number of days d. and (b) Use the equation you created in Part (a) to find the cost of renting a bike for 5 days.

OpenStudy (texaschic101):

C = 30 + 10d for 5 days.. C = 30 + 10(5) C = 30 + 50 C = 80 chart of values... C.....30.......40.......50.......60......70......80......90.....100 ------------------------------------------------- d.....0...........1........2........3..........4........5.......6.........7